OpenStudy (anonymous):

I'm going crazy over her. I've been to Khan academy and the website where I do my math but I can not for the life of me graph a quadratic. Can someone explain to me how to graph y=-4(x-6)^2+2

OpenStudy (charlotte123):

I use Wolframalpha ----> http://www.wolframalpha.com/input/?i=+y%3D-4%28x-6%29%5E2%2B2 <----

OpenStudy (amistre64):

i just make a U, or an upside down U

OpenStudy (charlotte123):

LOL! @amistre64

OpenStudy (anonymous):

That website is great. Thank you for telling me about it. But I also need to know how to solve it without a website.

OpenStudy (ddcamp):

Step 1: find the vertex If the equation is\[y=a(x-h)^2+k\] The vertex is at (h,k)

OpenStudy (charlotte123):

@Akjankins No problem!

OpenStudy (anonymous):

@DDCamp Alright, I'm following.

OpenStudy (amistre64):

y=-4(x-6)^2+2 ^ first term is negative, this models -x^2 --------------------------- let x = 0 to find the point it hits the y axis: y=-4(-6)^2+2 ------------------------------ y=-4(x-6)^2+2 ^ ^^ these represent the vertex ------------------------------- let y = 0 to define the points of intersection with the x axis 0 =-4(x-6)^2+2 , solve for x

OpenStudy (ddcamp):

Step 2: Find the direction (pretty easy) If a is positive, graph is U shaped. if a is negative, graph is upside down U shaped.

OpenStudy (ddcamp):

Step 3: Plot some points To get the best idea for the shape of the graph, plot points around the vertex (So if the vertex is at x=6, make a table for x= 4, 5, 7, & 8, then plot those points.)

OpenStudy (anonymous):

Thanks all of you.